# Irreducibility and periodicity in $\mathbb{Z}^{2}$ symbolic systems

### v1

@@ 1.1 Statement of the main result
Many mixing conditions have been introduced in the theory of $\mathbb{Z}^{d}$ symbolic dynamics, with the goal of finding a class of systems that behaves analogously to one-dimensional case. This line of investigation probably begins with Ruelle’s notion of specification [14], which was arose in his study of the theormodynamic formalism. Other aspects of the theory have lead to various other notions: block gluing, corner gluing, uniform filling and square mixing (e.g. [6, 13, 10, 12, 3, 7].
@@
Strongly irreducible systems behave in many respects like one-dimensional mixing shifts of finite type, and in $\mathbb{Z}^{2}$ the resemblence is even stronger. They also exhibit fewer logical complications than shifts of finite type. We summarize these similarities in the table below:
$\mathbb{Z}$-SFT
$\mathbb{Z}^{d}$-SFT
Strongly Irreducible
$\mathbb{Z}^{d}$-subshift
Strongly Irreducible
$\mathbb{Z}^{d}$-SFT
Language is
decidable
N/A (requires
finite description)
[8]
Unique measure of
maximal entropy
No, but fully
supported
No, but fully
supported [4]
Total transitivity
implies mixing
Automatic
Automatic
Factor onto lower
entropy full shifts
[9]
[6]
Subshifts that factors
into $X$ embed in $X$
[2]
[10, 11]
Periodic points
exist / are dense
THIS PAPER
✓ if $d\leq 2$, [10, 11]
if $d\geq 3$
In the present paper we consider the question, raised in [5], of existence of periodic points in SI systems (without assuming the SFT propery). Here, a periodic point means a point with finite orbit under the shift action. For $\mathbb{Z}$-subshifts, an elegant argument by Bertrand [1, Theorem 1] and, independently, Weiss [5, Theorem 5.1] shows that strongly irreducible $\mathbb{Z}$-subshifts admit dense periodic points (see [5]), and as we have noted, the same is true in $\mathbb{Z}^{2}$ if we add an assumption that the subshift is an SFT.
Joshua Frisch has pointed out to us that the question of existence of periodic points in SI $\mathbb{Z}^{2}$ subshifts is heuristically related to the question of their existence in SI SFTs in $\mathbb{Z}^{3}$. Indeed, if $X$ is a SI $\mathbb{Z}^{2}$-subshift, then for every $k$, the $\mathbb{Z}$-subshift consisting of restrictions of its configurations to ×$\mathbb{Z}\times\{1,\ldots,k\}$ (with $\mathbb{Z}$ acting by horizontal shift) is again SI, so by Weiss’s result it contains $\mathbb{Z}$-periodic points. Similarly, if $Y$ is a SI $\mathbb{Z}^{3}$ SFT, then for every $k$ one can consider locally admissible configurations on $\mathbb{Z}^{2}\times\{1,\ldots,k\}$, and this is a SI $\mathbb{Z}^{2}$ SFT, so by Lightwood’s result it contains $\mathbb{Z}^{2}$-periodic points. In both cases, there is no obvious way to extend this lower-dimensional periodicity to periodic points of the original system, but it lends some support to the idea that periodic points may exist.
@@ 1.2 Discussion
  • Syndeticity: Every closed $R_{n}$-ball contains $u,v$ for some $(u,v)\in\mathcal{W}_{n}$.
@@
This construction never yields a SI subshift in dimension $d=1$, for, by Bertrand/Weiss’s result, if $X_{(R_{n})}$ were SI it would contain periodic points, which it does not. Our main theorem shows that this argument cannot be applied in higher dimensions, but the following construction shows that for at least some sequences $R_{n}$, the subshift $X_{(R_{n})}$ is not SI.
Example 1.2.
For $A=\{0,1\}$ we shall construct a sequence $R_{n}\rightarrow\infty$ such that $X_{(R_{n})}\subseteq\{0,1\}^{\mathbb{Z}^{2}}$ is not SI. More precisely we construct an integer sequence $n_{k}\rightarrow\infty$ such that each integer $\ell$ divides some $n_{k}$, and a sequence $r_{k}\rightarrow\infty$. We then set $R_{n}=\min\{r_{k}\mid k\in\mathbb{N}\>,\>n|n_{k}\}$, and take $X_{(R_{n})}$ to be our example.
Fix a one-dimensional aperiodic sequence $z\in\{0,1\}^{\mathbb{N}}$ and an auxiliary sequence $s_{n}\rightarrow\infty$ with the property that every subword of length $s_{n}$ of $z$ contains an $n$-aperiodic pair. A square pattern will be said to be $z$-lined if the pattern along its sides (excluding the corners) are an initial segment of $z$. Every square $N\times N$ pattern can be extended to a $z$-lined pattern of dimensions $(N+2)\times(N+2)$. Observe that
If we tile a large square with $z$-lined $(2s_{n}+2)\times(2s_{n}+2)$ patterns, then every $2s_{n}\times 2s_{n}$ sub-pattern $b$ will contain the initial $s_{n}$-segment of $z$ along some row or column, a therefore $b$ contains an $n$-aperiodic pair.
Given a sequence $r_{1}<\ldots,<r_{k}$, we say that an $r_{k}\times r_{k}$ pattern is acceptable (with respect to $r_{1},\ldots,r_{k}$ and $n_{1},\ldots,n_{k}$) if for every $m\leq k$, every $r_{m}\times r_{m}$ subpattern of $a$ admits an $n_{m}$-aperiodic pair..
We shall inductively define integers $n_{1},n_{2},\ldots$ and $r_{1}<r_{2}<\ldots$ with $r_{m}\geq 2s_{n_{m}}$, and acceptable patterns $x^{(n)}$ over $\{0,1\}$, each $x^{(n)}$ being a square pattern centered at $z$ and extending $x^{(n+1)}$, as follows.
For $k=1$ let $n_{1}=1$ and let $x^{(1)}$ be any non-$1$-periodic pattern of dimension $2s_{1}\times 2s_{1}$ centered at 00 and with $x_{0}^{(1)}=0$. Set $r_{1}=2s_{1}$. Evidently $x^{(1)}$ is acceptable, since the only $r_{1}\times r_{1}$ sub-pattern of $x^{(1)}$ is $x^{(1)}$ itself.
Now suppose that $n_{1},\ldots,n_{k}$ and $r_{1}<\ldots<r_{k}$ have been defined with $r_{m}\geq 2s_{m}$, and let $x^{(k)}$ be an acceptable pattern with respect to them.
Choose the minimal $g\in\mathbb{N}$ and the lexicographically least pattern $b\in\{0,1\}^{[-g,g]^{2}}$ with the property that
$b_{0}=1$
and such that the $r_{k}\times r_{k}$ pattern $y^{(k)}$ given by
$y^{(k)}=\left\{\begin{array}[]{cc}y_{u}^{(k)}=x_{u}^{(k)}&u\in[r_{k}\times r_{k}]^{2}\setminus[-g,g]^{2}\\ b_{u}&u\in[-g,g]^{2}\end{array}\right.$
is also acceptable with respect to $r_{1}\leq\ldots\leq r_{k}$ and $n_{1}\leq\ldots\leq n_{k}$. Certainly such a $y^{(k)}$ exists, e.g. one is obtained by flipping all bits in $x^{(k)}$.
Let
$n_{k+1}=k(r_{k}+2)$
and let
$r_{k+1}\geq 2s_{n_{k}}$
be a multiple of $n_{k}$.
To define $x^{(k+1)}$, first extend $x^{(k)},y^{(k)}$ to $z$-lined patterns $\widehat{x}^{(k)},\widehat{y}^{(k)}$ of dimensions $(r_{k}+2)\times(r_{k}+2)$. Then extend $\widehat{x}^{(k)}$ to an $r_{k+1}\times r_{k+1}$ pattern by surrounding $\widehat{x}^{(k)}$ with translates of $\widehat{y}^{(k)}$ subject only to the requirement that the original $\widehat{x}^{(k)}$ not be in one of the corners of $x^{(k+1)}$.
We claim that $x^{(k+1)}$ is acceptable with respect to $r_{1},\ldots,r_{k+1}$ and $n_{1},\ldots,n_{k+1}$. Indeed,
  • The only $r_{k+1}\times r_{k+1}$ sub-pattern is $x^{(k+1)}$ itself, and $\widehat{x}_{0}^{(k)}=0\neq 1=\widehat{x}_{(0,r_{k+1})}^{(k+1)}$ (because $(0,r_{k+1})$ is the center of a copy of $\widehat{y}^{(k)}$).
  • For $1\leq m\leq k$, if $b$ is an $r_{m}\times r_{m}$ sub-pattern of $\widehat{x}^{(k+1)}$. then either it is a sub-pattern of $x^{(k)}$ or a translate of $y^{(k)}$, in which case $b$ contains an $n_{m}$-aperiodic pair because $x^{(k)},y^{(k)}$ are acceptable, or else $b$ intersects the boundary of a translate of $\widehat{x}^{(k)}$ or $\widehat{y}^{(k)}$ in a segment of length at least $r_{m}/2$, and there one sees an initial segment of $z$ of length $\geq r_{m}/2-1=s_{m}$, so there is an $n_{m}$-aperiodic pair.
Now, the sequence $x^{(k)}$ increases to $x\in\{0,1\}^{\mathbb{Z}^{2}}$ and clearly $x\in X_{(R_{n})}$ for $R_{n}$ defined as above. On the other hand, form the construction it is clear that for every gap $g$, if we set flip $x_{0}=1$, there is no way to redefine $x|_{[-g,g]^{2}\setminus\{0\}}$ in a manner that will keep $x$ in $X_{(R_{n})}$. For we have constructed the $x^{(k)}$ precisely so that however we fill in the symbols in $[-g,g]^{2}\setminus\{0\}$, this pattern will have been tried at some step in the construction, and inserting it results in a periodic $r_{k+1}\times r_{k+1}$ rectangle tiled by copies of $\widehat{y}^{(k)}$ and this pattern is $n_{k}=2r_{k}+1$ periodic.
@@
This construction demonstrates why controlling the effects of gluing is so difficult: when a finite region is cut out of a configuration and another pattern glued in its place, the change potentially can create periodicity at any one of infinitely many “scales”. Thus, if one wants to glue in a way that prevents periodicity, one potentially must satisfy infinitely many constraints, arising from symbols arbitrarily far away. At the same time, we are free to re-define only finitely many symbols near the boundary where the gluing occurred. So there are not always enough degrees of freedom to ensure aperiodicity.
@@
We note that the example above produced very special sequences $R_{n}$. We do not know if some sequences $(R_{n})$ can give rise to SI subshifts $S_{(R_{n})}$. Even if the answer is negative, it seems possible that subshifts defined as above can lead to SI aperiodic systems in the followins sense.. Given a rule $S$ for gluing configurations together, every subshift $X$ defines a strongly irreducible closure $\left\langle X\right\rangle_{S}=\bigcup_{n=1}^{\infty}X_{n}$, where $X_{1}=X$ and $X_{n+1}$ is obtained by gluing patterns from $X_{n}$ in all possible ways using $S$. One can then ask if there is a sequence $R_{n}>1$ and a gluing rule $S$ such that $\left\langle X_{(R_{n})}\right\rangle_{s}$ is aperiodic?
@@ 1.3 Ideas from the construction
Since we are free to re-define the symbols in $(E\cup F)^{c}$, we can try to resolve this by creating a new aperidic pair, e.g. by “moving” $u$ to the set $(E\cup F)^{c}$, whose size is at least proportional to the boundary of $E$. But this strategy faces two problems:
@@
    There may not be enough space in $(E\cup F)^{c}$ to create new $n$-aperiodic pairs for every pair that was destroyed. Indeed, we can only assume that $(E\cup F)^{c}$ is proportional to the length of the boundary of $E$, whereas, potentially, the number of sites belonging to $n$-aperiodic pairs in $E$ may be proportional to the size of $E$ itself. Actually the situation is not quite this bad, since we need not worry about pairs $(u,v)$ for which both $u$ and $v$ in the same set $E$ or $F$, so, assuming we bound the distance between $u,v$ by some $R_{n}$, we need “only” worry about $n$-aperiodic pairs that lie within $R_{n}$ of the boundary of $E$. Nevertheless, since $R_{n}$ grows with $n$, when $E$ is finite the number of pairs that need fixing easily exceeds the number of sites we can use to fix them.
@@
    When one member $u\in E$ of an aperiodic pair $(uv)\in E\times F$ is “moved” from $E$ to a site $u^{\prime}\in(E\cup F)^{c}$, we may not be able to ensure that $u=u^{\prime}\bmod n$, so, even though we may be able to ensure different symbols at $u^{\prime}$ and $v$, the pair $(u^{\prime},v)$ is no an $n$-aperiodic pair. Instead, we could try to pair $u^{\prime}$ with a new site $v^{\prime}$ to get an $n$-aperiodic pair, but it may happen that all points that agree with $u^{\prime}$ modulo $n$ are already part of an aperiodic pair, or that this cannot be done without violating the sparsity condition discussed in (1).
@@
  • Aperiodicity: Every $W\in\mathcal{W}_{n}$ contains a pair $u\neq v$ with $u=v\bmod n$ and $x_{n}\neq x_{v}$ Furthermore, the pattern $x|_{W}$ will be “robust” in the sense that its properties can be preserved despite local changes.
@@
  • Sparsity: For every finite region $E$, the set $\{u\in\mathbb{Z}^{2}\setminus E\mid d(u,E)>g\}$ is larger than $\sum_{n}|E\cap(\cup\mathcal{W}_{n})|$
@@
It is important to keep in mind that the structure $(\mathcal{W}_{n})$ and other auxiliary structure supporting it will not be encoded in the configurations of the subshift. This makes the structure flexible, and when gluing configurations these structure may change also far from the gluing boundary. We shall call this auxiliary structure a certificate. Thus, in order to belong to our subshift, a point must admit compatible certificate, and our task will be to show that, when two configurations $x,y$ are glued together to give $z$, we can simultaneously “glue” their certificates to give an admissible certificate for $z$.
1.4 Standing notation
Throughout the paper we write $a\gg b$ for “$a$ sufficiently large relative to $b$”. Thus, a statement of the form “If $a\gg b$, then…” means “For every $b$, if $a$ is sufficiently large in a manner depending on $b$, then…”. A statement such as “Since $a\gg b$ we have…” means “Since we are free to choose $a$ arbitrarily large relative ot $b$, we can choose it large enough that…”.
@@ 1.6 Acknowledgment
I would like to thank Joshua Frisch for bringing this problem to my attention.

### v2

@@ 1.1 Statement of the main result
Many mixing conditions have been introduced in the theory of $\mathbb{Z}^{d}$ symbolic dynamics, with the goal of finding a class of systems that behaves analogously to one-dimensional case. This line of investigation probably begins with Ruelle’s notion of specification [15], which was arose in his study of the thermodynamic formalism. Other aspects of the theory have lead to various other notions: block gluing, corner gluing, uniform filling and square mixing (e.g. [6, 14, 10, 13, 3, 7]).
@@
Strongly irreducible systems behave in many respects like one-dimensional mixing shifts of finite type, and in $\mathbb{Z}^{2}$ the resemblance is even stronger. They also exhibit fewer logical complications than shifts of finite type. We summarize these similarities in the table below:
$\mathbb{Z}$-SFT $\mathbb{Z}^{d}$-SFT Strongly Irreducible $\mathbb{Z}^{d}$-subshift Strongly Irreducible $\mathbb{Z}^{d}$-SFT
Language is decidable Undefined (requires recursive description) [8]
Mixing implies unique measure of maximal entropy No, but fully supported No, but fully supported [4]
Total transitivity implies mixing Automatic Automatic
Factor onto lower entropy full shifts [9] [6]
Subshifts that factors into $X$ embed in $X$ [2] [10, 11]
Periodic points exist / are dense THIS PAPER ✓ if $d\leq 2$, [10, 11] if $d\geq 3$
In the present paper we consider the question, raised in [5], of existence of periodic points in SI systems (without assuming the SFT property). Here, a periodic point means a point with finite orbit under the shift action. For $\mathbb{Z}$-subshifts, an elegant argument by Bertrand [1, Theorem 1] and, independently, Weiss [5, Theorem 5.1] shows that strongly irreducible $\mathbb{Z}$-subshifts admit dense periodic points (see [5]), and as we have noted, the same is true in $\mathbb{Z}^{2}$ if we add an assumption that the subshift is an SFT.
Joshua Frisch has pointed out to us that the question of existence of periodic points in SI $\mathbb{Z}^{2}$ subshifts is heuristically related to the question of their existence in SI SFTs in $\mathbb{Z}^{3}$. Indeed, if $X$ is a SI $\mathbb{Z}^{2}$-subshift, then for every $k$, the $\mathbb{Z}$-subshift consisting of restrictions of its configurations to $\mathbb{Z}\times\{1,\ldots,k\}$ (with $\mathbb{Z}$ acting by horizontal shift) is again SI, so by Weiss’s result it contains $\mathbb{Z}$-periodic points. Similarly, if $Y$ is a SI $\mathbb{Z}^{3}$ SFT, then for every $k$ one can consider locally admissible configurations on $\mathbb{Z}^{2}\times\{1,\ldots,k\}$, and this is a SI $\mathbb{Z}^{2}$ SFT, so by Lightwood’s result it contains $\mathbb{Z}^{2}$-periodic points. In both cases, there is no obvious way to extend this lower-dimensional periodicity to periodic points of the original system, but it lends some support to the idea that periodic points may exist.
@@ 1.2 Discussion
  • Syndeticity: Every closed $R_{n}\times R_{n}$-square contains $u,v$ for some $(u,v)\in\mathcal{W}_{n}$.
@@
Subshifts of the form $X_{(R_{n})}$ are never SI in dimension $d=1$, for, by Bertrand/Weiss’s result, if $X_{(R_{n})}$ were SI it would contain periodic points, which it does not. Our main theorem shows that this argument cannot be applied in higher dimensions, i.e. the non-existence of periodic points in $X_{(R_{n})}$ does not in itself rule out the possibility that $X_{(R_{n})}$ is SI. However, for at least some sequences $R_{n}$, we are able to show that the subshift $X_{(R_{n})}$ is not SI.
Exampls 1.2.
For $A=\{0,1\}$ there exists a sequence $R_{n}\rightarrow\infty$ such that $X_{(R_{n})}\subseteq\{0,1\}^{\mathbb{Z}^{2}}$ is not SI.
@@
The construction is given in Section 2. It demonstrates well why controlling the effects of gluing is so difficult: when a finite region is cut out of a configuration and another pattern glued in its place, the change potentially can create periodicity at any one of infinitely many “scales”. Thus, if one wants to glue in a way that prevents periodicity, one potentially must satisfy infinitely many constraints, arising from symbols arbitrarily far away. At the same time, we are free to re-define only finitely many symbols near the boundary where the gluing occurred. So there are not always enough degrees of freedom to ensure aperiodicity.
@@
We note that the example above produced very special sequences $R_{n}$. We do not know if some sequences $(R_{n})$ can give rise to SI subshifts $X_{(R_{n})}$. Even if the answer is negative, it seems possible that subshifts defined as above can lead to SI aperiodic systems in the following sense. Given a rule $S$ for gluing configurations together, every subshift $X$ defines a strongly irreducible closure $\left\langle X\right\rangle_{S}=\bigcup_{n=1}^{\infty}X_{n}$, where $X_{1}=X$ and $X_{n+1}$ is obtained by gluing patterns from $X_{n}$ in all possible ways using $S$. One can then ask if there is a sequence $R_{n}>1$ and a gluing rule $S$ such that $\left\langle X_{(R_{n})}\right\rangle_{s}$ is aperiodic?
@@ 1.3 Ideas from the construction
Since we are free to re-define the symbols in $(E\cup F)^{c}$, we can try to resolve this by creating a new aperiodic pair, e.g. by “moving” $u$ to the set $(E\cup F)^{c}$, whose size is at least proportional to the boundary of $E$. But this strategy faces two problems:
@@
    There may not be enough space in $(E\cup F)^{c}$ to create new $n$-aperiodic pairs for every pair that was destroyed. Indeed, we can only assume that $(E\cup F)^{c}$ is proportional to the length of the boundary of $E$, whereas, potentially, the number of sites belonging to $n$-aperiodic pairs in $E$ may be proportional to the size of $E$ itself. Actually the situation is not quite this bad, since we need not worry about pairs $(u,v)$ for which both $u$ and $v$ are in the same set $E$ or $F$, so, assuming we bound the distance between $u,v$ by some $R_{n}$, we need “only” worry about $n$-aperiodic pairs that lie within $R_{n}$ of the boundary of $E$. Nevertheless, since $R_{n}$ grows with $n$, when $E$ is finite the number of pairs that need fixing easily exceeds the number of sites we can use to fix them.
@@
    When one member $u\in E$ of an aperiodic pair $(u,v)\in E\times F$ is “moved” from $E$ to a site $u^{\prime}\in(E\cup F)^{c}$, we may not be able to ensure that $u=u^{\prime}\bmod n$, so, even though we may be able to ensure different symbols at $u^{\prime}$ and $v$, the pair $(u^{\prime},v)$ is not an $n$-aperiodic pair. Instead, we could try to pair $u^{\prime}$ with a new site $v^{\prime}$ to get an $n$-aperiodic pair, but it may happen that all points that agree with $u^{\prime}$ modulo $n$ are already part of an aperiodic pair, or that this cannot be done without violating the sparsity condition discussed in (1).
@@
  • Aperiodicity: Every $W\in\mathcal{W}_{n}$ contains a pair $u\neq v$ with $u=v\bmod n$ and $x_{u}\neq x_{v}$ Furthermore, the pattern $x|_{W}$ will be “robust” in the sense that its properties can be preserved despite local changes.
@@
  • Sparsity: For every finite region $E$, the set $\{u\in\mathbb{Z}^{2}\setminus E\mid d(u,E)\leq g\}$ is larger than $\sum_{n}|E\cap(\cup\mathcal{W}_{n})|$ .
@@
It is important to keep in mind that the structure $(\mathcal{W}_{n})$ and other auxiliary structure supporting it will not be encoded in the configurations of the subshift. This makes the structure flexible, and when gluing configurations these structures may change also far from the gluing boundary. We shall call this auxiliary structure a certificate. Thus, in order to belong to our subshift, a point must admit a compatible certificate, and our task will be to show that, when two configurations $x,y$ are glued together to give $z$, we can simultaneously “glue” their certificates to give an admissible certificate for $z$.
1.4 Standing notation: $a\gg b$
Throughout the paper we write $a\gg b$ for “$a$ sufficiently large relative to bb”. Thus, a statement of the form “If $a\gg b$, then…” means “For every $b$, if $a$ is sufficiently large in a manner depending on $b$, then…”. A statement such as “Since $a\gg b$ we have…” means “Since we are free to choose $a$ arbitrarily large relative to $b$, we can choose it large enough that…”.
@@ 1.6 Acknowledgment
I would like to thank Joshua Frisch for bringing this problem to my attention. I am also grateful to the anonymous referees for suggesting many corrections to the first version of the manuscript.
2 Example 1.2
Our goal is to show that there exists a sequence $R_{n}\rightarrow\infty$ such that $X_{(R_{n})}\subseteq\{0,1\}^{\mathbb{Z}^{2}}$ is not SI. The subshift $X_{(R_{n})}$ was defined in the Section 1.2.
The framework for the construction is as follows. We will define an integer sequence $n_{k}\rightarrow\infty$ such that every integer $\ell$ divides some $n_{k}$, and another integer sequence $r_{k}\rightarrow\infty$, with $r_{k}$ a multiple of $n_{k}$. We say that an $r_{k}\times r_{k}$ pattern is acceptable (with respect to $r_{1},\ldots,r_{k}$ and $n_{1},\ldots,n_{k}$, whose mention we shall usually leave implicit) if for every $m\leq k$, every $r_{m}\times r_{m}$ subpattern of $a$ admits an $n_{m}$-aperiodic pair. Set
$R_{n}=\min\{r_{k}\mid k\in\mathbb{N}\>,\>n|n_{k}\}$
We then take $X_{(R_{n})}$ as be our example.
Claim 2.1.
If $x\in\{0,1\}^{2}$ contains an $r_{k}\times r_{k}$ subpattern $b$ that is not acceptable, then $x\notin X_{(R_{n})}$.
Proof.
Indeed, let $1\leq m\leq k$ be an index such that $b$ does not contain an $n_{m}$-periodic pair. Then $R_{n_{m}}=r_{m}$, because $n_{m^{\prime}}<n_{m}$ for $m^{\prime}<m$, and hence $n_{m}$ cannot divide any $n_{m^{\prime}}$ with $m^{\prime}<m$, and hence $b$ is a pattern on an $R_{n_{m}}\times R_{n_{m}}$-ball that does not contain any $n_{m}$-aperiodic pairs, showing that $x\notin X_{(R_{n})}$. ∎
We now turn to the construction proper. Fix a one-dimensional aperiodic sequence $z\in\{0,1\}^{\mathbb{N}}$ and an auxiliary sequence $s_{n}\rightarrow\infty$ with the property that every subword of length $s_{n}$ of $z$ contains an $n$-aperiodic pair. A square pattern will be said to be $z$-lined if the pattern along each of its sides, read in the counter-clockwise direction and excluding the corners, is an initial segment of $z$. Every square $N\times N$ pattern can be extended to a $z$-lined pattern of dimensions $(N+2)\times(N+2)$.
We now define the sequences $n_{1}<n_{2}<\ldots$ and $r_{1}<r_{2}<\ldots$ inductively, satisfying $r_{m}\geq 2s_{n_{m}}+2$, together with $r_{k}\times r_{k}$-patterns $x^{(k)}$ that are centered at the origin $0=(0,0)$, are acceptable in the sense defined above, and such that $x^{(k)}$ extends $x^{(k-1)}$. The definition is as follows:
Let $n_{1}=1$ and $r_{1}=2s_{1}+3$. Let $x^{(1)}$ be any centered $r_{1}\times r_{1}$-pattern with the symbol zero at the origin (i.e., $x_{0}^{(1)}=0$) and also containing the symbol $1$ (so it is $1$-aperiodic). Trivially, $x^{(1)}$ is acceptable, since the only $r_{1}\times r_{1}$ sub-pattern of $x^{(1)}$ is $x^{(1)}$ itself and we chose it to be $1$-aperiodic.
Suppose that $n_{1}<\ldots<n_{k}$, $r_{1}<\ldots<r_{k}$, and $x^{(1)},\ldots,x^{(k)}$ have been defined as stated. Choose the minimal $g_{k}\in\mathbb{N}$ and, given $g_{k}$, the lexicographically least pattern $b^{(k)}\in\{0,1\}^{[-g_{k},g_{k}]^{2}}$, such that the pair $(g_{k},b^{(k)})$ satisfy
$b_{0}^{(k)}=1$
and such that the $r_{k}\times r_{k}$ pattern $y^{(k)}$ given by
$y^{(k)}=\left\{\begin{array}[]{cc}y_{u}^{(k)}=x_{u}^{(k)}&u\notin[-g_{k},g_{k}]^{2}\\ b_{u}^{(k)}&u\in[-g_{k},g_{k}]^{2}\end{array}\right$.
is acceptable. To justify taking the least such pair note that such pairs $(g_{k},b^{(k)})$ exists, e.g. one is obtained by taking $g_{k}=\frac{1}{2}(r_{k}-1)$ and flipping all bits in $x^{(k)}$, so $b_{u}^{(k)}=1-x_{u}^{(k)}$.
Let
$n_{k+1}=k(r_{k}+2)$
(in particular, $k|n_{k+1}$), and let
$r_{k+1}\geq 2s_{n_{k}}+2$
be any odd multiple of $n_{k+1}$.
To define $x^{(k+1)}$, first extend $x^{(k)},y^{(k)}$ to $z$-lined patterns $\widehat{x}^{(k)},\widehat{y}^{(k)}$ of dimensions $(r_{k}+2)\times(r_{k}+2)$. Then extend $\widehat{x}^{(k)}$ to a centered $r_{k+1}\times r_{k+1}$ pattern by surrounding $\widehat{x}^{(k)}$ with translates of $\widehat{y}^{(k)}$.
Claim 2.2.
$x^{(k+1)}$ is acceptable.
Proof.
Just observe that
  • The only $r_{k+1}\times r_{k+1}$ sub-pattern is $x^{(k+1)}$ itself, and $\widehat{x}_{0}^{(k)}=0\neq 1=\widehat{x}_{(0,n_{k+1})}^{(k+1)}$ (because $(0,n_{k+1})$ is the center of a copy of $\widehat{y}^{(k)}$). So $0=(0,0)$ and $(0,n_{k+1})$ are an $n_{k}$-aperiodic pair.
    • For $1\leq m\leq k$, if $b$ is an $r_{m}\times r_{m}$ sub-pattern of $\widehat{x}^{(k+1)}$, then either it is a sub-pattern of $x^{(k)}$ or of (a translate of) $y^{(k)}$, in which case $b$ contains an $n_{m}$-aperiodic pair because $x^{(k)},y^{(k)}$ are acceptable, or else $b$ intersects the boundary of $\widehat{x}^{(k)}$ or of (a translate of) $\widehat{y}^{(k)}$ in a segment of $z$ of length at least $r_{m}/2$, in which we are guaranteed, be definition of $z$ and $s_{m}$, to see a initial segment of $z$ of length $\geq r_{m}/2-1=s_{m}$, which contains an $n_{m}$-aperiodic pair.
Claim 2.3.
If we replace the central $g_{k},g_{k}]^{2}$ pattern in $x^{(k+1)}$ with $b^{(k)}$ (equivalently, replace the central copy of $x^{(k)}$ with $y^{(k)}$), then the resulting pattern in no longer acceptable.
Proof.
The resulting pattern is tiled by copies of $\widehat{y}^{(k)}$, whose dimensions are $n_{k+1}\times n_{k+1}$, and hence contains no $n_{k+1}$-aperiodic pair. ∎
Corollary 2.4.
The sequence $(g_{k},b^{(k)})$ is strictly increasing in the lexicographic order on pairs $(g,b)$ with $g\in\mathbb{N}$ and $b\in\{0,1\}^{[-g,g]2}.$ In particular, $g_{k}\rightarrow\infty$.
Proof.
The previous claim shows that the sequence $b^{(k+1)}\neq b^{(k)}$, and in fact, since the central $r_{m}\times r_{m}$ pattern in $x^{(k)}$ is $x^{(m)}$, it shows that $b^{(k)}\neq b^{(m)}$ for all $m<k$. Thus $(g_{k},b^{(k)})$ does not contain any repetitions. The fact that it is strictly increasing now follows from our choice of the minimal candidate at each stage. ∎
Since $x^{(k+1)}$ extends $x^{(k)}$ and these are centered square patterns of increasing side length, these patterns increases to a configuration $x\in\{0,1\}^{\mathbb{Z}^{2}}$.
Claim 2.5.
$x\in X_{(R_{n})}$. In particular, $X_{(R_{n})}\neq\emptyset$.
Proof.
Given $n$, by definition there exists $k$ with $R_{n}=r_{k}$ and $n|n_{k}$. Every $R_{n}\times R_{n}$-subpattern $a$ of $x$ is a sub-pattern of some $x^{(m)}$ with $m>k$, and since $x^{(m)}$ is acceptable, $a$ must contain an $n_{k}$-aperiodic pair, and hence an $n$-aperiodic pair (since $n|n_{k}$). ∎
Claim 2.6.
$X_{(R_{n})}$ is not SI .
Proof.
Fix $g\in\mathbb{N}$ and let us show that $X_{(R_{n})}$ is not SI with gap $g$. Specifically, we claim that if we set $x_{0}=1$, then there is no way to redefine $x$ on $g,g]^{2}\setminus\{0\}$ in a manner that the resulting configuration $x^{\prime}$ belongs to $X_{(R_{n})}$.
Indeed, let $b\in\{0,1\}^{[-g,g]^{2}}$ with $b_{0}=1$ and let $x^{\prime}$ be as above. Then either $b\neq b^{(k)}$ for all $k$, which by definition means that the central $r_{k}$-pattern in $x^{\prime}$ is not acceptable for any $k$ with $g_{k}>g$, or else $b=b^{(k)}$ for some $k$, in which case by Claim 2.3, the central $r_{k+1}\times r_{k+1}$-pattern in $x^{\prime}$ is not acceptable. In both cases $x^{\prime}\notin X_{(R_{n})}$. ∎
This concludes the construction.